A 500.0Hz sound source is moving away from an observer and approaching a wall at a speed of 20m/s. What is the frequency distinguished by the observer of the sound waves reaching him after reflecting on the wall? (Speed of sound on air: 335 m/s)
1
Expert's answer
2019-12-04T08:38:43-0500

The frequency emitted by the moving object:


"f_1=\frac{335+0}{335-20}500=531.746\ Hz"

The frequency distinguished by the observer of the sound waves reaching him after reflecting on the wall:


"f_2=\frac{335+20}{335-0}531.746=563.5\ Hz"